# From percolation transition to Anderson localization in one-dimensional speckle potentials

### v1

@@ 4.1 Transfer-matrix simulations
To evaluate the mean transmission $\langle T\rangle$ in the semi-classical regime, we first performed numerical simulations based on the 1D transfer-matrix method. We discretize the stationary Schrödinger equation $\hbar^{2}\partial_{x}^{2}\psi(x)/(2m)+V(x)\psi(x)=E\psi(x)$ with a red-detuned potential $V(x)$ over a system of length $L$. Expressing the wave function $\psi=\{\psi_{n}\}_{0\leq n<N}$ and the potential $V={V_{n}}$ on $N=L/a$ lattice sites, we obtain the effective tight-binding model:
@@
Fig. 3(a) first shows the average transmission as a function of the system size $L$, computed numerically for $\hbar_{\text{eff}}=0.01$ and for energies in the vicinity of the percolation threshold. The transmission exhibits a marked exponential decay. Fits of the data to $\propto\exp[-L/\chi(E)]$ allow us to extract the localization length $\chi$ as a function of energy, which is plotted in Fig. 3(b) for several values of $\hbar_{\text{eff}}$ closer and closer to the classical limit. Remarkably, as $\hbar_{\text{eff}}\to 0$, the localization length converges to the classical result (10) (long-dashed curve) that characterizes the percolation transition. The figure also reveals how this classical phase transition gradually disappears for finite $\hbar_{\text{eff}}$: the function $\chi(E)$ smoothly evolves from a strict algebraic divergence at $E=V_{0}$ when $\hbar_{\text{eff}}=0$ to an ultra-steep yet continuous behavior at this point in the semi-classical $\hbar_{\text{eff}}\ll 1$. This is the first important result of the paper.
@@ 4.2 Semi-classical theory
To explain the numerical results of Fig. 3(b), we now present a theoretical calculation of the localization length in the semi-classical regime. To this end, we first notice that for energies $E\simeq V_{0}$ in the regime $\hbar_{\text{eff}}\ll 1$, particle propagation is still expected to be mainly governed by independent transmission events $T_{b}$ through individual barriers:
@@
like in the classical limit [see Eq. (8)]. From a quantum-mechanical perspective, the approximation (19) amounts to neglecting multiple Fabry–Pérot interferences between successive barriers, which, in contrast, become dominant in the deep quantum regime $\hbar_{\text{eff}}\gg 1$ of Anderson localization. Under this approximation, the average transmission in the semi-classical limit is given by
$\langle T\rangle\simeq\langle T_{b}\rangle^{N_{b}}\simeq\exp[-L/\chi(E)],\ \ \chi^{-1}(E)\equiv-\rho\,\text{ln}\langle T_{b}\rangle,$ (20)
@@
$\chi^{-1}(E)=-\rho\,\text{ln}\left\{1-\int_{0}^{\infty}d\omega\int_{-\infty}^{V_{0}}dV\,\frac{P(V,\omega)}{1+\exp\left[\frac{2\pi}{\hbar\omega}(E-V)\right]}\right\}.$ (22)
@@ 4.3.2 Localization length near the threshold
where $\alpha=(1-\sqrt{2})\zeta(1/2)\Gamma(7/4)/\sqrt{\pi}$ and $\beta=2\sqrt{\pi}(1-2\sqrt{2})\zeta(-1/2)\Gamma(5/4)$. At leading order in the semi-classical approach, the localization length thus diverges as $\smash{\hbar_{\text{eff}}^{-1/2}}$ at the percolation threshold. The approximation (26) is shown as a dotted curve in Fig. 3(b).
@@ 4.3.3 Localization length below the threshold
We finally turn our attention to the most interesting regime: the behavior of localization below the threshold. To obtain an approximate expression for $\chi$ in this case, we use the asymptotic expansion $\text{Li}_{1/2}(-x)\simeq-2\sqrt{\text{ln}(x)/\pi}+\pi^{3/2}/(12\,\text{ln}^{3/2}x)$ when $x\to\infty$ in Eq. (23). In terms of the inverse localization length, this leads to:
@@ 5.1 Classical limit
To further characterize transport across the percolation transition in the semi-classical regime, we also examined the statistical distribution of the transmission, $P(T)$, for small values of $\hbar_{\text{eff}}$. In the classical limit $\hbar_{\text{eff}}=0$, a straightforward calculation presented in the appendix yields
@@ 5.2 Numerical simulations at finite $\hbar_{\text{eff}}$
We now turn to the transmission distribution at finite $\hbar_{\text{eff}}$. Like for the localization length, we expect a smoothing of $P(T)$ across the percolation threshold. Transfer-matrix simulations of $P(T)$ shown in Fig. 4 for three energies around $E=V_{0}$ confirm this behavior:
@@
As the threshold is crossed at finite $\hbar_{\text{eff}}$, $P(T)$ evolves continously from a bimodal form for $E<V_{0}$ toward a distribution sharply peaked at $T=1$ for $E\geq V_{0}$. The most intriguing phenomenon here occurs below the threshold, where the classical doubled-delta law (29) is replaced, in the semi-classical regime, by a much smoother bimodal distribution. Remarkably, this bimodal structure is not captured by conventional 1D localization theory for uncorrelated Gaussian disorder. To illustrate this failure, we reproduce in Fig. 5 the numerically obtained bimodal distribution for $E<E_{0}$, and compare it with the exact solution of Abrikosov’s Fokker-Planck (or 1D DMPK) equation for the transmission distribution derived for Gaussian and uncorrelated disorder [42],
$P(T)=\frac{2}{T^{2}\sqrt{\pi}(L/\xi)^{3/2}}\int_{\text{acosh}(1/\sqrt{T})}^{\infty}\frac{x\exp(-x^{2}\xi/L-L/4\xi)}{\sqrt{\cosh^{2}x-1/T}}dx.$ (30)
@@
Figure 5: Transmission distribution $P(T)$ below the percolation threshold ($E-V_{0}=-0.005V_{0}$), in the semi-classical regime with $\hbar_{\text{eff}}=0.005$. A comparison between the numerical results and the DMPK prediction (30), using either $\xi$ or $\chi/4$ as the localization length (dotted and dashed curves), highlights the inadequacy of this approach. In contrast, the comparison with the bimodal law (41) (solid green curve) shows excellent agreement. For reference, we also display $P(T)$ in the deep quantum regime ($\hbar_{\text{eff}}=1$), for which the DMPK equation becomes accurate again. All simulations are performed for a system size $L=100\sigma$ and involve binning over $10^{5}$ disorder realizations.
For reference, we also show in Fig. 5 a numerical calculation of $P(T)$ in the quantum regime $\hbar_{\text{eff}}=1$. In this case, $P(T)$ takes the more usual log-normal shape and Eq. (30) becomes accurate again. Like for the localization length, we attribute the breakdown of Eq. (30) in the semi-classical regime to the strong influence of spatial correlations and non-Gaussian statistics of the red-detuned disorder near the percolation threshold. In the quantum regime, tunnelling through barriers dominates, making the disorder correlations less relevant and restoring the applicability of the DMPK theory developed for uncorrelated Gaussian potentials.
Turning back to the semi-classical regime, the shape of $P(T)$ strongly resembles the bimodal distribution characteristic of diffusive transport in multi-channel waveguides [48]. It is important to stress, however, that the bimodality here a priori has a fundamentally different origin. In quasi-1D waveguides, diffusion indeed stems from the presence of many transverse channels $N$; the localization length $\chi\sim N\ell\gg\ell$ is much longer than the mean free path $\ell$, so a diffusive window appears for system sizes $\ell\ll L\ll\chi$. In contrast, the red-detuned speckle potential near the percolation threshold is strictly 1D ($N=1$), so diffusion cannot be attributed to a multi-channel effect. Instead, we again trace its origin to disorder correlations: for $\hbar_{\text{eff}}\ll 1$ the scattering mean free path presumably becomes much smaller than $\chi$, such that the 1D system can support diffusive transport. Understanding how this bimodal distribution emerges from a semi-classical description is the objective of the next section.
@@ 5.3 The bimodal law from semi-classical theory
$P(T)=\frac{1}{T}\int_{-\infty}^{\infty}dt\,\exp\big[\!-it\ln T+N_{b}\ln\varphi(t)\big],$ (32)
@@
$\varphi(t)\equiv\langle e^{it\ln T_{j}}\rangle=\int_{0}^{\infty}d\omega\int_{-\infty}^{V_{0}}dV\,P(V,\omega)\exp\Big[\!-it\ln\Big(1+e^{-\frac{2\pi}{\hbar\omega}(E-V)}\Big)\Big],$ (35)
@@
$\varphi(t)\simeq\int_{0}^{\infty}d\omega\int_{-\infty}^{E}dV\,P(V,\omega)\exp\Big[\!-ite^{-\frac{2\pi}{\hbar\omega}(E-V)}\Big],$ (36)
@@
This result exactly coincides with the bimodal law characteristics of diffusive systems [48], which we here recover from a semi-classical approach in the vicinity of the percolation threshold. This is the second important result of the paper. Equation (41) is shown in Fig. 5, and describes very well the transfer-matrix simulations, even up to rather small transmission values.
@@ 5.4 Quasi-ballistic regime
where $x=1/\cosh^{2}x$ and taking $t=L/\chi$. This relation leads to a reasonable agreement of the numerical results, at least for not too small $T$ values.
@@ 6 From semi-classical to deep quantum regime
So far, we have focused on the semi-classical regime $\hbar_{\text{eff}}\ll 1$. In this final section, we investigate how the localization length behaves as one crosses over into the deep quantum regime $\hbar_{\text{eff}}\gg 1$. To characterize this crossover, one may consider either the localization length defined from the average transmission, $\chi=-L/\ln\langle T\rangle$, or the one defined from the average logarithmic transmission, $\xi=-L/\langle\ln T\rangle$. As shown in the previous sections, $\chi$ is the quantity that properly interpolates to the critical cluster size in the limit $\hbar_{\text{eff}}\to 0$. However, it is also known that $\chi$ is not an appropriate measure of localization in the quantum regime, due to strong transmission fluctuations. In that regime, the relevant quantity is $\xi$, since it is constructed from $\langle\ln T\rangle$ which is self-averaging. Put differently, when $\hbar_{\text{eff}}\ll 1$, transport is dominated by free propagation within clusters and is thus physically captured by $\chi$ (see Sec. 3), whereas in the quantum regime $\hbar_{\text{eff}}\gg 1$ transport is dominated by tunneling which is more faithfully captured by $\xi$. To describe the crossover between these two opposite regimes, we make the—admittedly somewhat arbitrary—choice to focus on $\xi$, as it is well defined in the deep quantum limit while still exhibiting, like $\chi$, a divergence at the percolation threshold in the semiclassical regime [Eq. (16)].
@@ 6.2 Perturbation theory in the deep quantum regime
$\frac{1}{\xi(E)}\simeq\frac{1}{\sigma}\frac{E_{\sigma}}{E}\;\frac{1}{\hbar_{\text{eff}}^{4}}\;\int_{-\infty}^{0}\mathrm{d}u\;\exp\Big(-\frac{u^{2}}{2}\Big)\cos\left(2u\sqrt{2{E}/{E_{\sigma}}}\right)+\mathcal{O}(\hbar_{\text{eff}}^{-6}).$ (43)
@@ 7 Conclusion
In this work, we investigated both numerically and theoretically the classical-to-quantum crossover between the percolation transition and Anderson localization in a red-detuned speckle potential. This crossover is controlled by an effective Planck constant, $\hbar_{\text{eff}}$, defined as the ratio between the de Broglie wavelength and the disorder correlation length. We found that as the system departs from the classical limit $\hbar_{\text{eff}}=0$, the characteristic algebraic divergence of the cluster length $\chi$ at the percolation threshold progressively evolves into a continuous localization length. For small but finite $\hbar_{\text{eff}}$, the correlated and non-Gaussian nature of the speckle potential plays a crucial role, causing the standard DMPK description of Anderson localization in uncorrelated disorder to break down. Using a semi-classical approach, we developed a consistent theoretical description of the localization length in this regime, correctly interpolating with the classical limit. Just below the percolation threshold, we also identified the emergence of a bimodal transmission law typical of diffusive transport—usually absent in 1D models with uncorrelated Gaussian disorder. We further provided an analytical derivation of this law within our semi-classical framework. Finally, we explored the global behavior of the localization length as $\hbar_{\text{eff}}$ is varied over several orders of magnitude, up to the deep quantum regime $\hbar_{\text{eff}}\gg 1$ where standard quantum perturbation theory becomes applicable.
@@
Our results provide an example of a situation in which Anderson localization exhibits non-universal properties due to the correlated nature of the disorder, as already emphasized, e.g., in Ref. [50, 51]. In the present case, properly accounting for these properties is essential to accurately capture the quantum-to-classical crossover. Our study also illustrates an interesting crossover between a regime featuring a genuine phase transition and another regime where no such transition exists. In disordered systems, similar situations arise, for instance, in dimensional crossovers from $d=3$ to $d<3$, or in crossovers between different symmetry classes such as in two-dimensional systems transitioning between the unitary and quantum-Hall classes [52], or between the unitary and orthogonal classes in three dimensions [53] where two distinct Anderson transitions occur.
@@
To deepen our understanding of the connection between percolation and Anderson localization, it would be worthwhile to investigate whether a generalized 1D DMPK framework can be derived for non-Gaussian correlated disorder, capturing the change of the transmission distribution from log-normal to bimodal observed in the present work. This might be done, for instance, using the recent approach of [54]. It would also be highly valuable to explore the three-dimensional case, where a classical percolation transition exists while an Anderson transition occurs deep in the quantum regime. How these two transitions emerge or disappear as $\hbar_{\text{eff}}$ is increased remains unclear. Finally, probing the dynamics of Anderson localization could provide an original perspective on the percolation–Anderson crossover, for instance through the expansion of wave packets or the motion of their center-of-mass via the quantum boomerang effect [55, 56].
@@
This work was supported by the project Localization of Waves of the Simons Foundation (Grant No. 1027116, M.F.).

### v2

@@ 4.1 Transfer-matrix simulations
To evaluate the mean transmission $\langle{T}\rangle$} in the semi-classical regime, we first perform numerical simulations based on the 1D transfer-matrix method. We discretize the stationary Schrödinger equation $\hbar^{2}\partial_{x}^{2}\psi(x)/(2m)+V(x)\psi(x)=E\psi(x)$ with a red-detuned potential $V(x)$ over a system of length $L$. Expressing the wave function $\psi=\{\psi_{n}\}_{0\leq n<N}$ and the potential $V={V_{n}}$ on $N=L/a$ lattice sites, we obtain the effective tight-binding model:
@@
Figure 3(a) first shows the average transmission as a function of the system size $L$, computed numerically for $\hbar_{\text{eff}}=0.01$ and for energies in the vicinity of the percolation threshold. The transmission exhibits a marked exponential decay. Fits of the data to $\propto\exp[-L/\chi(E)]$ allow us to extract the localization length $\chi$ as a function of energy, which is plotted in Fig. 3(b) for several values of $\hbar_{\text{eff}}$ closer and closer to the classical limit. Remarkably, as $\hbar_{\text{eff}}\to 0$, the localization length converges to the classical result (10) (long-dashed curve) that characterizes the percolation transition. The figure also reveals how this classical phase transition disappears for finite $\hbar_{\text{eff}}$: the function $\chi(E)$ smoothly evolves from a strict algebraic divergence at $E=V_{0}$ when $\hbar_{\text{eff}}=0$, to an ultra-steep yet continuous behavior at this point as soon as $\hbar_{\text{eff}}$ is nonzero. This is the first important result of the paper.
@@ 4.2 Semi-classical theory
To explain the numerical results of Fig. 3(b), we now present a theoretical calculation of the localization length in the semi-classical regime. To this end, we first note that in the deep quantum regime $\hbar_{\text{eff}}\gg 1$ of Anderson localization, statistical correlations between transmission events typically arise from the accumulation of phase factors associated with multiple quantum-mechanical reflections between successive barriers. For energies $E\simeq V_{0}$ in the semi-classical regime $\hbar_{\text{eff}}\ll 1$, in contrast, such reflections are negligible because the transmission through each barrier is very close to unity. As a result, transmission events $T_{b}$ through individual barriers are essentially independent, and
@@
like in the classical limit [see Eq. (8)]. Under this approximation, the average transmission in the semi-classical regime is given by
${\langle{T}\rangle}\simeq{\langle{T_{b}}\rangle}^{N_{b}}\simeq\exp[-L/\chi(E)],\ \ \chi^{-1}(E)\equiv-\rho\,\ln{\langle{T_{b}}\rangle},$ (20)
@@
$\chi^{-1}(E)=-\rho\,\text{ln}\left\{1-\int_{0}^{\infty}d\omega\int_{-\infty}^{V_{0}}dV\,\frac{P(V,\omega)}{1+\exp[\frac{2\pi}{\hbar\omega}(E-V)]}\right\}.$ (22)
@@ 4.3.2 Localization length near the threshold
where $\alpha=(1-\sqrt{2})\zeta(1/2)\Gamma(7/4)/\sqrt{\pi}$ and $\beta=2\sqrt{\pi}(1-2\sqrt{2})\zeta(-1/2)\Gamma(5/4)$. The localization length thus diverges as $\smash{\hbar_{\text{eff}}^{-1/2}}$ at the percolation threshold. The approximation (26) is shown as a dotted curve in Fig. 3(b).
@@ 4.3.3 Localization length below the threshold
We finally turn our attention to the most interesting regime: the behavior of localization below the threshold. To obtain an approximate expression for $\chi$ in this case, we use the asymptotic expansion $\text{Li}_{1/2}(-x)\simeq-2\sqrt{\ln(x)/\pi}+\pi^{3/2}/(12\,\ln^{3/2}x)$ when $x\to\infty$ in Eq. (23). In terms of the inverse localization length, this leads to:
@@ 5.1 Classical limit
To further characterize transport across the percolation transition in the semi-classical regime, let us also examine the statistical distribution of the transmission, $P(T)$, for small values of $\hbar_{\text{eff}}$. In the classical limit $\hbar_{\text{eff}}=0$, a straightforward calculation presented in the appendix yields
@@ 5.2 Numerical simulations at finite $\hbar_{\text{eff}}$
We now turn to the transmission distribution at finite $\hbar_{\text{eff}}$. Like for the localization length, we expect a smoothing of $P(T)$ across the percolation threshold. Transfer-matrix simulations of $P(T)$ shown in Fig. 4 for three energies around $E=V_{0}$ confirm this statement:
@@
As the threshold is crossed for finite $\hbar_{\text{eff}}$, $P(T)$ evolves continuously from a bimodal form for $E<V_{0}$ toward a distribution sharply peaked at $T=1$ for $E\geq V_{0}$. The most intriguing phenomenon here occurs below the threshold, where the classical doubled-delta law (29) is replaced, in the semi-classical regime, by a much smoother bimodal distribution. Remarkably, this bimodal structure is not captured by conventional 1D localization theory for uncorrelated Gaussian disorder. To illustrate this failure, we reproduce in Fig. 5 the numerically obtained bimodal distribution for $E<E_{0}$, and compare it with the exact solution of Abrikosov’s Fokker-Planck (or 1D DMPK) equation for the transmission distribution derived for Gaussian and uncorrelated disorder [42],
$P(T)=\frac{2}{T^{2}\sqrt{\pi}(L/\xi)^{3/2}}\int_{\text{acosh}(1/\sqrt{T})}^{\infty}\frac{x\exp(-x^{2}\xi/L-L/4\xi)}{\sqrt{\cosh^{2}x-1/T}}\,{\rm d}x.$ (30)
@@
Figure 5: Transmission distribution $P(T)$ below the percolation threshold ($E-V_{0}=-0.005V_{0}$), in the semi-classical regime with $\hbar_{\text{eff}}=0.005$. A comparison between the numerical results and the DMPK prediction (30), using either $\xi$ or $\chi/4$ as the localization length (dotted and dashed curves), highlights the inadequacy of this approach. In contrast, the comparison with the bimodal law (41) (solid green curve) shows excellent agreement. For reference, we also display $P(T)$ in the deep quantum regime ($\hbar_{\text{eff}}=1$), for which the DMPK equation becomes accurate again (solid red curve). All simulations are performed for a system size $L=100\sigma$ and involve binning over $10^{5}$ disorder realizations.
For reference, we also show in Fig. 5 a numerical calculation of $P(T)$ in the quantum regime $\hbar_{\text{eff}}=1$. In this case, $P(T)$ takes the more usual log-normal shape and Eq. (30) becomes accurate again. Like for the localization length, we attribute the breakdown of Eq. (30) in the semi-classical regime to the strong influence of spatial correlations and non-Gaussian statistics of the red-detuned disorder near the percolation threshold. In the quantum regime, tunneling through barriers dominates, making the disorder correlations less relevant and restoring the applicability of the DMPK theory developed for uncorrelated Gaussian potentials.
Turning back to the semi-classical regime, the shape of $P(T)$ strongly resembles the bimodal distribution characteristic of diffusive transport in multi-channel waveguides [48]. It is important to stress, however, that the bimodality here has a fundamentally different origin a priori. In quasi-1D waveguides, diffusion indeed stems from the presence of many transverse channels $N$; the localization length $\xi\sim N\ell\gg\ell$ is much longer than the mean free path $\ell$, so a diffusive window appears for system sizes $\ell\ll L\ll\xi$. In the present work, however, the random potential is strictly 1D ($N=1$) so diffusion cannot be attributed to a multi-channel effect. Understanding how this bimodal distribution emerges from a semi-classical description and what is the diffusive mechanism behind it is the objective of the next section.
@@ 5.3 The bimodal law from semi-classical theory
$P(T)=\frac{1}{T}\int_{-\infty}^{\infty}{\rm d}t\,\exp\big[\!-it\ln T+N_{b}\ln\varphi(t)\big],$ (32)
@@
$\varphi(t)\equiv{\langle{e^{it\ln T_{j}}}\rangle}=\int_{0}^{\infty}d\omega\int_{-\infty}^{V_{0}}dV\,P(V,\omega)\exp[\!-it\ln(1+e^{-\frac{2\pi}{\hbar\omega}(E-V)})],$ (35)
@@
$\varphi(t)\simeq\int_{0}^{\infty}d\omega\int_{-\infty}^{E}dV\,P(V,\omega)\exp[\!-ite^{-\frac{2\pi}{\hbar\omega}(E-V)}],$ (36)
@@
This result exactly coincides with the bimodal law characteristic of diffusive systems [48], which we here recover from a semi-classical approach in the vicinity of the percolation threshold. This is the second important result of the paper. Equation (41) is shown in Fig. 5, and describes very well the transfer-matrix simulations, even up to rather small transmission values.
Physically, we conjecture that the bimodal distribution (41) is associated with a diffusive process in which the particle of energy $E$ is multiply scattered by barriers whose height is close to $E$. For such barriers, the transmission coefficient is broadly distributed around 1/2 due to quantum reflection, while the peaks at $T=0$ and $T=1$ are associated with events in which the particle is either purely reflected or purely transmitted. Within this picture, the characteristic length $\chi$ plays a role analogous to a mean free path.
@@ 5.4 Quasi-ballistic regime
where $T=1/\cosh^{2}x$ and taking $t=L/\chi$. This relation leads to a reasonable agreement of the numerical results, at least for not too small $T$ values.
@@ 6 From semi-classical to deep quantum regime
So far, we have focused on the semi-classical regime $\hbar_{\text{eff}}\ll 1$. In this final section, we investigate how the localization length behaves as one crosses over into the deep quantum regime $\hbar_{\text{eff}}\gg 1$. To characterize this crossover, one may consider either the localization length defined from the average transmission, $\chi=-L/\ln{\langle{T}\rangle}$, or the one defined from the average logarithmic transmission, $\xi=-L/{\langle{\ln T}\rangle}$. As shown in the previous sections, $\chi$ is the quantity that properly interpolates to the critical cluster size in the limit $\hbar_{\text{eff}}\to 0$. However, it is also known that $\chi$ is not an appropriate measure of localization in the quantum regime, due to strong transmission fluctuations. In that regime, the relevant quantity is $\xi$, since it is constructed from $\langle\ln T\rangle$ which is self-averaging. Put differently, when $\hbar_{\text{eff}}\ll 1$, transport is dominated by free propagation within clusters and is thus physically captured by $\chi$ (see Sec. 3), whereas in the quantum regime $\hbar_{\text{eff}}\gg 1$ transport is dominated by tunneling which is more faithfully captured by $\xi$. To describe the crossover between these two opposite regimes, we make the—admittedly somewhat arbitrary—choice to focus on $\xi$, as it is well defined in the deep quantum limit while still exhibiting, like $\chi$, a divergence at the percolation threshold in the semi-classical regime [Eq. (16)].
@@ 6.2 Perturbation theory in the deep quantum regime
$\frac{1}{\xi(E)}\simeq\frac{1}{\sigma}\frac{E_{\sigma}}{E}\;\frac{1}{\hbar_{\text{eff}}^{4}}\;\int_{-\infty}^{0}\mathrm{d}u\;\exp(-\frac{u^{2}}{2})\cos\left(2u\sqrt{2{E}/{E_{\sigma}}}\right)+\mathcal{O}(\hbar_{\text{eff}}^{-6}).$ (43)
@@
For the sake of clarity, we summarize in Table 1 the definitions, physical interpretations and regime of relevance of the two localization lengths $\chi$ and $\xi$ introduced in the manuscript.
Quantity Definition Physical meaning Relevant regime
$\chi$ $\rho\,\ln{\langle{T}\rangle}]^{-1}$ Mean size of classically Semi-classical regime
allowed regions $\hbar_{\mathrm{eff}}\ll 1$
$\xi$ $\rho\,{\langle{\ln T_{b}}\rangle}]^{-1}$ Exponential decay length Quantum regime
of quantum eigenstates $\hbar_{\mathrm{eff}}\gg 1$
Table 1: Summary of the definitions, physical interpretations, and regime of relevance of the two localization lengths introduced in the manuscript. $T_{b}$ is the transmission through a single barrier and $\rho$ is the density of barriers.
@@ 7 Conclusion
In this work, we investigated both numerically and theoretically the classical-to-quantum crossover between the percolation transition and Anderson localization in a red-detuned speckle potential. This crossover is controlled by an effective Planck constant, $\hbar_{\text{eff}}$, defined as the ratio between the de Broglie wavelength and the disorder correlation length. We found that, as the system departs from the classical limit $\hbar_{\text{eff}}=0$, the characteristic algebraic divergence of the cluster length $\chi$ at the percolation threshold progressively evolves into a continuous localization length. For small but finite $\hbar_{\text{eff}}$, the correlated and non-Gaussian nature of the speckle potential plays a crucial role, causing the standard DMPK description of Anderson localization in uncorrelated disorder to break down. Using a semi-classical approach, we developed a consistent theoretical description of the localization length in this regime, correctly interpolating with the classical limit. Below the percolation threshold, we also identified the emergence of a bimodal transmission distribution, usually absent in 1D models with uncorrelated Gaussian disorder, and which we associate with a diffusive motion of the particle quantum-mechanically transmitted and reflected near the potential maxima. We further provided an analytical derivation of this law within our semi-classical framework. Finally, we explored the global behavior of the localization length as $\hbar_{\text{eff}}$ is varied over several orders of magnitude, up to the deep quantum regime $\hbar_{\text{eff}}\gg 1$ where standard quantum perturbation theory becomes applicable.
Although Anderson localization was originally formulated for diagonal disorder with short-range correlations, its scope has been extended over the years to more general types of disorder. Our results provide an example of a situation in which Anderson localization in non-Gaussian, correlated disorder exhibits original yet non-universal properties, as already emphasized, e.g., in Refs. [50, 51]. In the present case, properly accounting for the disorder statistics is essential to accurately capture the quantum-to-classical crossover. In particular, the correlated nature of the potential is crucial for the semi-classical regime (4) to be well defined, even though we expect our results in the vicinity of the percolation threshold—most notably the bimodal distribution—to be generic and largely independent of the details of the short-range disorder correlation function owing to the universal character of the transition. In the 1D geometry considered here, the bounded nature of the potential is also essential for the existence of a percolation threshold, but this would not be the case in higher dimension.
@@
Our study also illustrates an interesting crossover between a regime featuring a genuine phase transition and another regime where no such transition exists. In disordered systems, similar situations arise, for instance, in dimensional crossovers from $d=3$ to $d<3$, or in crossovers between different symmetry classes such as in two-dimensional systems transitioning between the unitary and quantum-Hall classes [52], or between the unitary and orthogonal classes in three dimensions [53] where two distinct Anderson transitions occur.
The semi-classical regime explored in this work should be experimentally accessible in cold-atom experiments with optical speckle potentials, where the effective Planck constant can be tuned by adjusting the disorder amplitude. While earlier experiments [9, 10, 11, 12] were deliberately designed to avoid classical trapping, with $\hbar_{\text{eff}}$ typically ranging from $0.5$ to $6$, the limit $\hbar_{\text{eff}}\ll 1$ should be achievable by employing larger disorder amplitudes. Moreover, experiments now offer the possibility to precisely control the energy of atoms loaded into disordered potentials [58], opening the way to a precise characterization of the critical behavior of $\chi(E)$.
@@
To deepen our understanding of the connection between percolation and Anderson localization, it would be worthwhile to investigate whether a generalized 1D DMPK framework can be derived for non-Gaussian correlated disorder, capturing the change of the transmission distribution from log-normal to bimodal observed in the present work. This might be done, for instance, using the approach of Ref. [54]. It would also be highly valuable to explore the three-dimensional case, where a classical percolation transition exists while an Anderson transition occurs deep in the quantum regime. How these two transitions emerge or disappear as $\hbar_{\text{eff}}$ is increased remains unclear. Finally, probing the dynamics of Anderson localization could provide an original perspective on the percolation–Anderson crossover, for instance through the expansion of wave packets or the motion of their center-of-mass via the quantum boomerang effect [55, 56].
@@
This work was supported by the project Localization of Waves of the Simons Foundation (Grant No. 1027116, M.F.), and by Agence Nationale de la Recherche under Grant No ANR-23-PETQ-0001 Dyn1D France 2030.